# Binaural Unmasking in Practical Use: Perceived Level of Phase-inverted Speech in Environmental Noise

### v1

@@ Abstract
We aim to develop a technology that makes the sound from earphones and headphones easier to hear without increasing the sound pressure or eliminating ambient noise. To this end, we focus on harnessing the phenomenon of binaural unmasking through phase reversal in one ear. Specifically, we conduct experiments to evaluate the improvement of audibility caused by the phenomenon, using conditions that approximate practical scenarios. We use speech sounds by various speakers, including women, and noises that can be encountered in daily life (urban environmental sounds, cheers) to verify the effects of binaural unmasking under conditions close to practical situations. The results of experiments using the Japanese language showed that (i) speech in a noisy environment is perceived to be up to about 6 dB louder with phase reversal in one ear, and (ii) a certain effect (improvement of audibility by 5 dB or more) is obtained for all speakers and noises targeted in this study. These findings demonstrate the effectiveness of binaural unmasking attributed to interaural phase differences in practical scenarios.
@@ 1 Introduction
However, most of the speech that has been studied so far has been read by a single male [8, 9, 10, 11], and there is little research involving various speakers with different formant patterns, including women. In other words, the effect on a variety of speakers is not clear. Moreover, the noises that have been studied are mostly mechanically generated ones such as broad-band Gaussian noise [8], speech spectrum noise [9], and speech weighted noise [10]. Therefore, the effect on “natural” noises found in everyday life (such as urban environmental sounds) is also not clear.
@@
  • Investigation of the effect for a variety of speakers by using various speakers with different formant patterns, including women.
@@ 2.2.1 Binaural Unmasking
When listening to a 500-Hz pure tone in a noise identical at the ears, inverting the phase of the target sound in one ear can lower the detection threshold by about 12–15 dB [19]. However, this effect declines to about 3 dB as the frequency of the target sound increases, and stabilizes at about 3 dB for all higher frequencies, up to at least 4 kHz [20].
@@ 3.1 Purpose and Overview
In the experiment, participants listened through headphones to an original signal (Original: $\rm{S_{0}}$) and a signal in anti-phasic conditions (Comparison: variations of $\rm{S_{\pi}}$), both played with an identical noise at both ears ($\rm{N_{0}}$). The initial gain of the Original signal (i.e., the gain before the volume adjustment) had been determined with the intention of creating an initial state of “audible but difficult to hear” through the process described in Appendix G. The participants were asked to compare the two sounds and adjust the volume of the Original until it was perceived as having the same level as the Comparison. In most cases, the Comparison was more audible than the Original, which means the volume of the Original needed to be increased to match the Comparison. If the Comparison was less audible, the volume of the Original was decreased. By comparing the two signal levels before and after the adjustment, we measured how much easier (or more difficult) it became to hear the speech in noise due to the phase reversal of the speech in one ear. Details will be described later in Section 3.3.
This experiment involved human subjects. Approval of all ethical and experimental procedures and protocols was granted by Sony Bioethics Committee (no. 25-R-0011).
@@ 3.2 Participants
The experiment was conducted with 17 participants; however, one participant was excluded for the reason described later in Section 4. Therefore, the experimental results were compiled based on data from 16 participants. The 16 participants were recruited from the corporate group that includes the company where the authors work (excluding the authors themselves). Among them, eleven were male and five were female, with two in their 20s, seven in their 30s, four in their 40s, and three in their 50s. All participants were native speakers of Japanese. Participants were limited to those who had no history of hearing issues in past hearing tests. However, note that we did not ask participants to submit the results of hearing tests; they only self-reported that they had never experienced hearing issues. No compensation or other benefits were provided to the participants, and the decision to participate was left to their free will. Each participant took part in the experiment only once.
@@ 3.3.1 Selection of Speech Signals
[…]
@@
[…]
@@ 3.3.3 Presentation of Signals/Noises and Volume Adjustment by Participants
Participants compare the two by alternately pressing the “Sound A” and “Sound B” buttons and then use the spin button to adjust the volume of the speech in Sound A in 1-dB increments so that the perceived level of the speech in Sounds A and B feels the same. Note that only the volume of the speech is adjusted, and the volume of the noise remains unchanged. Once the comparison and volume adjustment are completed, the participant presses the “Next” button. This action records the adjusted volume and loads the next stimuli (i.e., Sounds A and B). Simultaneously, the number next to “No.” automatically increments.
@@ 3.4.1 Devices Used for Sound Presentation
Fig. 4 shows the layout of the devices used for the experiment. The user interface described in 3.3.3 was run on a Windows PC. The participants used Sony headphones MDR-1AM2888https://electronics.sony.com/audio/headphones/all-headphones/p/mdr1am2-b to listen to the sounds. A Sony WALKMAN NW-A55999https://www.sony.com/electronics/support/digital-music-players-nw-nwz-a-series/nw-a55 was connected between the PC and the headphones, and the WALKMAN’s USB DAC (Digital to Analog Converter) function was utilized to set the volume displayed on the WALKMAN to 52. This corresponds to approximately 56.6 dB SPL [26]. The connection between the headphones and the WALKMAN was made using the original cable that comes with the headphones. The signal was output from PC to USB DAC AMP through the Audio Driver (ASIO).
@@ 4 Results
The results for the Dummy condition are shown in Table I, where the values represent the BHLD obtained from three to four Dummy trials conducted by each participant [^10] Each session included one or two dummy trials, and since each participant conducted three sessions, the total was three to four trials. As mentioned in Section 3.3.3, participants actually conducted four sessions each, but the first session was for practice and is not included in the results tally.. Note that P9 was excluded from the aggregation of experimental results due to the large volume adjustment. According to these results, each participant, except for P9, understood the purpose of the experiment and did not make substantial volume adjustments in conditions where there was no difference in audibility; in other words, they made correct judgments. Although participant P1 made a 4-dB adjustment once, the other two trials were correctly maintained at 0 dB, so it was deemed not to be an issue.
@@
In the aggregated results for all conditions (Fig. 6), there was no statistically significant difference between the BHLD for All Booster and Low Booster at $F_{c}$ = 1000 Hz ($p=0.75$, Welch’s t-test). Both BHLD values were about 6 dB. Specifically, the sample means for All Booster and Low Booster at $F_{c}$ = 1000 Hz were 5.91 and 5.80, respectively. The 95% confidence intervals (CIs) based on t-distribution (df = 53) were 5.42 to 6.37 and 5.27 to 6.30 for All Booster and Low Booster at $F_{c}$ = 1000 Hz, respectively, indicating that it is highly likely that the true population means fall within these ranges. For Low Booster, the BHLD increased as $F_{c}$ increased. On the other hand, for High Booster, the BHLD decreased as $F_{c}$ increased.
@@
For all signals and noises, a BHLD of more than 5 dB was obtained. For All Booster of Signals A, B, and C, the sample means were 5.91, 6.19, and 5.62, respectively, and the 95% CIs based on t-distribution (df = 17) were 4.96 to 6.76, 5.37 to 6.93, and 4.62 to 6.52, respectively. For All Booster of Noises A, B, and C, the sample means were 6.09, 4.98, and 6.58, respectively, and the 95% CIs based on t-distribution (df = 17) were 5.24 to 6.86, 5.74 to 7.35, and 4.11 to 5.77, respectively. As for Noise B, Low Booster at $F_{c}$ = 1000 Hz showed the largest BHLD (5.41), and the 95% CI based on t-distribution (df = 17) was 4.33 to 6.37.
@@ 5.3 Values and Concerns for Practical Applications
Considering practical use, it is necessary to reduce the discomfort in auditory perception. Based on the discussion in 5.2, we expect that by limiting the frequency band of phase inversion in one ear, a more effective BHLD can be achieved while avoiding the discomfort in sound localization caused by inverting the entire frequency range.
@@ 6 Conclusion
In this study, we investigated the effects of binaural unmasking through various anti-phasic conditions, focusing on variations in speakers and noises, with the aim of applying the phenomenon to engineering applications. The results of experiments targeting the Japanese language confirmed that anti-phasic presentation can make a speech, which was barely audible in a noisy environment, be perceived as up to about 6 dB louder. Improving the perceived level by 6 dB, or in other words, reducing the actual sound pressure by 6 dB, is equivalent to being able to extend the permissible safe listening time by 15 hours per week under a certain sound pressure condition. Although the degree of effect obtained varied depending on the target signals and the noises, a consistent effect (BHLD of at least 5 dB) was achieved for all target signals and noises addressed in this study.

### v2

@@ Abstract
We aim to develop a technology that makes the sound from earphones and headphones easier to hear without increasing the sound pressure or eliminating ambient noise. To this end, we focus on harnessing the phenomenon of binaural unmasking through phase reversal in one ear. Specifically, we conduct experiments to evaluate the improvement of audibility caused by the phenomenon, using conditions that approximate practical scenarios. We use speech sounds by various speakers and noises that can be encountered in daily life (urban environmental sounds, cheers) to verify the effects of binaural unmasking under conditions close to practical situations. The results of experiments using the Japanese language showed that (i) speech in a noisy environment is perceived to be up to about 6 dB louder with phase reversal in one ear, and (ii) a certain effect (improvement of audibility by 5 dB or more) is obtained for all speakers and noises targeted in this study. These findings demonstrate the effectiveness of binaural unmasking attributed to interaural phase differences in practical scenarios.
@@ 1 Introduction
However, most of the speech that has been studied so far has been read by a single male [8, 9, 10, 11], and there is little research involving various speakers with different formant patterns, including women. In other words, the effect on a variety of speakers is not clear. Moreover, the noises that have been studied are mostly mechanically generated ones such as broad-band Gaussian noise [8], speech spectrum noise [9], and speech weighted noise [10]. Therefore, the effect on noises found in everyday life (such as urban environmental sounds) is also not clear.
@@
  • Investigation of the effect for a variety of speakers by using speakers with different formant patterns, including women.
@@ 2.2.1 Binaural Unmasking
When listening to a 500-Hz pure tone in a noise identical at the ears, inverting the phase of the target sound in one ear can lower the detection threshold by about 12–15 dB [19]. However, this effect declines to about 3 dB as the frequency of the target sound increases, and stabilizes at about 3 dB for all higher frequencies, up to at least 4 kHz [20]. The effect of binaural unmasking becomes more pronounced as the masking noise level increases, that is, as the signal-to-noise ratio (SNR) decreases, and it diminishes as the SNR increases [6, 10, 21, 22, 23].
@@ 3.1 Purpose and Overview
In the experiment, participants listened through headphones to an original signal (Original: $\rm{S_{0}}$) and a signal in anti-phasic conditions (Comparison: variations of $\rm{S_{\pi}}$), both played with an identical noise at both ears ($\rm{N_{0}}$). The initial gain of the Original signal (i.e., the gain before the volume adjustment) had been determined with the intention of creating an initial state of “audible but difficult to hear” through the process described in Appendix G. The participants were asked to compare the two sounds and adjust the volume of the Original until it was perceived as having the same level as the Comparison. In most cases, the Comparison was more audible than the Original, which means the volume of the Original needed to be increased to match the Comparison. If the Comparison was less audible, the volume of the Original was decreased. By comparing the two signal levels before and after the adjustment, we measured the change in perceived level of the speech in noise resulting from the phase reversal of the speech in one ear. Details will be described later in Section 3.3.
This experiment involved human participants. Approval of all ethical and experimental procedures and protocols was granted by Sony Bioethics Committee (no. 25-R-0011).
@@ 3.2 Participants
The experiment was conducted with 17 participants; however, one participant was excluded for the reason described later in Section 4. Therefore, the experimental results were compiled based on data from 16 participants. The 16 participants were recruited from a corporate group that includes the authors’ company, excluding the authors themselves. Among them, eleven were male and five were female, with two in their 20s, seven in their 30s, four in their 40s, and three in their 50s. All participants were native speakers of Japanese. Participants were limited to those who had no history of hearing issues in past hearing tests. However, note that we did not ask participants to submit the results of hearing tests; they only self-reported that they had never experienced hearing issues. No compensation or other benefits were provided to the participants, and the decision to participate was left to their free will. Each participant took part in the experiment only once.
@@ 3.3.1 Selection of Speech Signals
The content of the speech used in the experiment was a sentence in Japanese from the ITA corpus [^1] https://github.com/mmorise/ita-corpus, specifically RECITATION324_180 “日本へ行くには船か飛行機が必要です,” which means “To go to Japan, you need a ship or an airplane.” The ITA corpus is a collection of sentences selected and modified for readability and public decency from literary works whose copyrights have expired [28]. RECITATION324_180 was chosen because it meets the following criteria, with the aim of selecting a sentence that would not distract the participants’ concentration:
@@
The speech signals utilized in the experiment were the recordings of RECITATION324_180 spoken by narrators or voice actors (one male, two females). Specifically, we used male/02_fast and female/02_fast from SpeedSpeech-JA-2022222https://ast-astrec.nict.go.jp/release/speedspeech_ja_2022/, and the “normal” version from Amitaro’s Voice Material Studio [^3] https://amitaro.net/. Hereafter, we refer to them as Signals A, B, and C, respectively. Signal A is a male voice, while Signals B and C are female voices with different formant patterns. The fact that two out of the three speakers were female was a result of using three speakers with different formant patterns. The spectrograms of Signals A, B, and C are shown in Fig. 9. It can be observed here that the voice of the second female speaker (Signal C) contains stronger high-frequency components. The duration of Signals A, B, and C is approximately 3 seconds each. The “fast” versions from SpeedSpeech-JA-2022 were chosen because the “normal” speaking rate felt slower than what is typically heard in everyday life. The details of how Signals A, B, and C were processed are described in Appendix F. Licensing information for the sound sources is listed in Table II of the Appendix.
@@ 3.3.3 Presentation of Signals/Noises and Volume Adjustment by Participants
Participants compare the two sounds by alternately pressing the “Sound A” and “Sound B” buttons. Then, they use the spin button to adjust the volume of the speech in Sound A in 1-dB increments until the perceived level of the speech in Sounds A and B feels the same. To prevent bias in participants’ responses, the user interface displayed volume as unlabeled numeric values, and participants were not informed that each increment represented 1 dB. Note that only the volume of the speech is adjusted, and the volume of the noise remains unchanged. Once the comparison and volume adjustment are completed, the participant presses the “Next” button. This action records the adjusted volume and loads the next stimuli (i.e., Sounds A and B). Simultaneously, the number next to “No.” automatically increments.
@@ 3.4.1 Devices Used for Sound Presentation
Fig. 4 shows the layout of the devices used for the experiment. The user interface described in 3.3.3 was run on a Windows PC. The participants used Sony headphones MDR-1AM2888https://electronics.sony.com/audio/headphones/all-headphones/p/mdr1am2-b to listen to the sounds. A Sony WALKMAN NW-A55999https://www.sony.com/electronics/support/digital-music-players-nw-nwz-a-series/nw-a55 was connected between the PC and the headphones, and the WALKMAN’s USB DAC (Digital to Analog Converter) function was utilized to set the volume displayed on the WALKMAN to 52. The connection between the headphones and the WALKMAN was made using the original cable that comes with the headphones. The signal was output from PC to USB DAC AMP through the Audio Driver (ASIO). Under this device configuration, the A-weighted equivalent continuous sound pressure level (LAeq) remained below 65 dBA, even when the speech volume was adjusted using the user interface. This LAeq is well below levels known to pose a risk to human hearing[29, 30].
@@ 4 Results
The results for the Dummy condition (where sound source A and B were identical) are shown in Table I. The values shown represent the BHLD obtained from three to four Dummy trials conducted by each participant [^10] Each session included one or two dummy trials, and since each participant conducted three sessions, the total was three to four trials. As mentioned in Section 3.3.3, participants actually conducted four sessions each, but the first session was for practice and is not included in the results tally.. Note that P9 was excluded from the aggregation of experimental results because the volume adjustment deviated from zero (greater than 4 dB). According to these results, each participant, except for P9, understood the purpose of the experiment and did not make substantial volume adjustments in conditions where there was no difference in audibility. Although participant P1 made a 4-dB adjustment once, the other two trials were maintained at 0 dB, so it was deemed not to be an issue.
@@
In the aggregated results for all conditions (Fig. 6), there was no statistically significant difference between the BHLD for All Booster and Low Booster at $F_{c}$ = 1000 Hz ($p=0.75$, Welch’s t-test). Both BHLD values were about 6 dB. Specifically, the sample means for All Booster and Low Booster at $F_{c}$ = 1000 Hz were 5.91 dB and 5.80 dB, respectively. The 95% confidence intervals (CIs) based on t-distribution (df = 53) were 5.42 dB to 6.37 dB and 5.27 dB to 6.30 dB for All Booster and Low Booster at $F_{c}$ = 1000 Hz, respectively, indicating that it is highly likely that the true population means fall within these ranges. For Low Booster, the BHLD increased as $F_{c}$ increased. On the other hand, for High Booster, the BHLD decreased as $F_{c}$ increased.
@@
For all signals and noises, a BHLD of more than 5 dB was obtained. For All Booster of Signals A, B, and C, the sample means were 5.91 dB, 6.19 dB, and 5.62 dB, respectively, and the 95% CIs based on t-distribution (df = 17) were 4.96 dB to 6.76 dB, 5.37 dB to 6.93 dB, and 4.62 dB to 6.52 dB, respectively. For All Booster of Noises A, B, and C, the sample means were 6.09 dB, 4.98 dB, and 6.58 dB, respectively, and the 95% CIs based on t-distribution (df = 17) were 5.24 dB to 6.86 dB, 5.74 dB to 7.35 dB, and 4.11 dB to 5.77 dB, respectively. As for Noise B, Low Booster at $F_{c}$ = 1000 Hz showed the largest BHLD (5.41 dB), and the 95% CI based on t-distribution (df = 17) was 4.33 dB to 6.37 dB.
@@ 5.3 Values and Concerns for Practical Applications
Considering practical use, it is necessary to reduce the discomfort in auditory perception. Based on the discussion in 5.2, we expect that by limiting the frequency band of phase reversal in one ear, a more effective BHLD can be achieved while avoiding the discomfort in sound localization caused by inverting the entire frequency range.
@@ 6 Conclusion
In this study, we investigated the effects of binaural unmasking through various anti-phasic conditions, focusing on variations in speakers and noises, with the aim of applying the phenomenon to engineering applications. The experimental results targeting the Japanese language confirmed that anti-phasic presentation can make a speech, which was barely audible in a noisy environment, be perceived as up to about 6 dB louder. An improvement of 6 dB in perceived level implies that, after phase reversal in one ear, the sound pressure can be reduced by 6 dB while maintaining the same perceived loudness as before the phase reversal. Reducing the sound pressure level by 6 dB increases the permissible safe listening time by about 15 hours per week under a given sound pressure condition. Although the degree of effect obtained varied depending on the target signals and the noises, a consistent effect (BHLD of at least 5 dB) was achieved for all target signals and noises addressed in this study.
@@ Appendix F Process for Aligning Volume of Speech Signals
Appendix G Determination of Initial Gains of Signals
Considering that listening difficulties in everyday life typically occur when the SNR is negative, conditions with positive SNR were excluded from the experiment. Under positive SNR, the target sound in our stimulus set remained clearly perceptible and was never fully masked by noise.
Furthermore, even when the SNR value was identical across stimulus sets, the degree of listening difficulty varied depending on the combination of target sound and noise. For this reason, the initial gain was adjusted based on human perceptual judgment.
The initial gain settings were determined by the authors, who did not participate in the experiment. Although the authors’ prior knowledge of the target phrases might have facilitated speech perception in noise, this does not reduce the listening difficulty for participants without such knowledge. Therefore, this methodological aspect was considered not to affect the experimental conditions.
The SNRs after setting the initial gains of the Original signals are shown in Table VI. In all conditions, the SNRs exhibited negative values.
@@ Appendix G Determination of Initial Gains of Signals
TABLE VI: SNRs of the stimuli after initial gain adjustment.
Signal Noise SNR [dB]
Signal A Noise A 25.1-25.1
Signal A Noise B 28.8-28.8
Signal A Noise C 19.1-19.1
Signal B Noise A 26.1-26.1
Signal B Noise B 31.1-31.1
Signal B Noise C 22.7-22.7
Signal C Noise A 25.1-25.1
Signal C Noise B 24.6-24.6
Signal C Noise C 23.5-23.5